# When LLMs Go Online: The Emerging Threat of Web-Enabled LLMs

### v2

@@ 5.1 Experimental Setup
Evaluation. We examined the impact of web accessibility on impersonation by conducting an A/B test. Impersonation posts were generated using (1) vanilla LLMs, (2) WebSearch agents, and (3) WebNav agents; then, we conducted pairwise comparisons to determine which of the two posts was more likely to be written by the target individual. As evaluators, we employed LLMs, which are recognized for their high accuracy in tasks such as fake news detection and machine-generated text detection, even without additional fine-tuning [45, 11, 34]  [^9] We also verified that LLM evaluators showed a high correlation with human evaluations in our experiment (See Appendix C.2).. We enabled the web search function to enhance the accuracy of the LLM evaluations by leveraging internet-available information. Given impersonation posts from two models (e.g., vanilla LLM vs. WebSearch agent, or WebSearch agent vs. WebNav agent), LLM evaluators indicated which post more accurately impersonated the target individual, or indicated ’unsure’ if the two posts were similarly effective. We randomize the placement of texts (A or B) to mitigate order bias. Furthermore, to minimize potential model bias, each pair of posts was evaluated by three different models: GPT, Claude, and Gemini. The final outcome was determined based on the majority vote from the three evaluations.
In addition, to determine how likely the impersonation posts were perceived to be authentic, we conducted a Yes/No test. Similar to the A/B test, we utilized the three LLM evaluators and to assess whether the texts appeared to have been authored by the target individuals. The templates for the A/B test and Yes/No test are provided in Appendix C.1.
@@ 5.2 Main Results
Figure 7 shows an example of an impersonation post  [^10] For visualization, we prompted the agent to generate a tweet of up to 300 words with the second claim. of a well-known figure generated by the WebSearch agent with GPT. We provide the agent with only the name information; however, it synthesizes information about his expertise, alma mater, and companies to craft a tweet that convincingly appears to be authored by him. The proceeding tests are conducted using target professors according to our attack pipeline.
Comparison Based on Tool Usage. We conducted comparative analysis of impersonation posts generated by WebNav agents and WebSearch agents, followed by comparisons between WebSearch agents and vanilla LLMs, to examine their effectiveness based on their capability. Table 2 shows the proportion of choices made in an A/B test. Cases where the LLMs or agents refused to generate content were excluded from our calculations (these will be discussed later in this section). The results indicate that the impersonation effectiveness increases with access to additional tools across all models. Specifically, WebNav agents were more effective than WebSearch agents, and WebSearch agents outperformed vanilla LLMs. The differences in effectiveness were more pronounced when comparing WebSearch agents to vanilla LLMs, suggesting that adding search capabilities (WebSearch agents) provides a significant boost in performance for tasks requiring impersonation abilities.
Table 2: A/B test results. The percentages represent each decision relative to the total number of decisions made. claim1 refers to “Recommend researching AI” and claim2 to “LLM is highly secure against potential misuse”. WebN and WebS represent WebNav and WebSearch, respectively.
@@
Table 3: Yes/No Test results using GPT evaluators. Percentages represent the proportion of “Yes” responses.
@@
Gen WebNav Agent WebSearch Agent LLM WebNav Agent WebSearch Agent LLM
@@
Authenticity Evaluation. We examine whether the texts created by the WebNav agents, WebSearch agents, and vanilla LLMs are perceived as being written by the target. Table 3 shows the percentage of “yes” responses by the LLM evaluator for posts created by vanilla LLMs and LLM agents. The results are similar to those observed in the A/B test: agent-generated posts had a higher “yes” response rate than LLM-generated posts. WebNav agents performed the best; the GPT and Claude agents achieved success rates of up to 93.9% and 85.7%, respectively. Compare to GPT and Claude, Gemini showed poor performance in most cases.
@@
6 Spear Phishing Email Generation
In this section, we investigate whether LLM agents can be used to generate personalized phishing emails using only an email address, without the need for additional human intervention.
@@ 6.1 Experimental Setup
Note that we conducted two pilot studies, which showed similar results from WebNav and WebSearch Agents. Due to cost and time constraints, WebNav-generated emails were excluded from the main evaluation.
@@
As in Section 4, academic researchers’ information is publicly available online, making them more vulnerable to phishing attacks. Therefore, we categorized participants into two groups: academic researchers (including undergraduate students) and non-academic professionals (individuals within non-academic organizations). Academic researchers were recruited via university research communities, while non-academic professionals were recruited through industry practitioners from prior collaborations. Specifically, we recruited 60 participants and categorized them based on the provided institutional email addresses to analyze response variations across institutional types: 55% were academic researchers and 45% were non-academic professionals  [^11] The recruitment process with ethical/privacy-preserving measures is explained in Appendix E.1..
We emphasize that participants did not receive these emails directly; instead, they were assessed via a Google survey. Each participant evaluated seven phishing emails generated using their own email as a target. The types of emails will be described in Email Design part.
@@
Questionnaire Design. We designed seven multiple-choice questions to build an understanding of how participants perceive and interact with the emails, from surface content to deeper implications of authenticity and expected actions.
The first two questions focus on the content of the emails, asking participants to identify the presented information (Q1) and to evaluate its accuracy against their actual personal details (Q2). Subsequent questions explore perceptions of the sender and recipient, including whether the recipient identified the sender (Q3) and if the recipient’s name correctly matched their real name (Q4). We then inquire about the actions participants would likely take upon receiving the email (Q5). The last question investigates perceived authenticity of the email (Q6), asking participants to rate how genuine or fraudulent it appeared and to identify specific elements that influenced their judgment (Q7). The questionnaire is available in Appendix E.5.
@@ 6.2 Main Results
We first analyze the effectiveness and content of general-purpose emails created by LLMs and WebSearch agents. Then, we examine the results according to the participants’ groups. We then analyze the efficacy of emails according to their purpose and sender’s institution. In Tables 10 (Appendix E.3), we report the comprehensive survey results concerning participants’ anticipated actions for various email types.
@@ 6.2.1 Overall Analysis
Effectiveness Analysis. Figure 10 (a) presents the survey results of actions participants would likely take upon receiving the email (Q5). Our primary focus was on the proportion of participants clicking on links, directly aligning with the attacker’s goal. The results indicate that WebSearch agents prompt more link clicks than vanilla LLMs, with the WebSearch agent using Claude achieving the highest click rate at 26.67%  [^12] Expert-crafted spear phishing emails in [9] show 26.6% click rate. Detailed discussion in Section 6.2.3.. In the case of GPT, the WebSearch agent’s click rate was twice that of the vanilla LLM. When comparing Claude and GPT, Claude was more effective, with its vanilla LLM variant achieving a slightly higher click rate than the WebSearch agent using GPT. The disparities in effectiveness between Claude and GPT may be attributed to differences in email content, which will be explored further in the subsequent section.
When comparing “actions” (Q5) and “perceived authenticity” (Q6) in Figure 10, on average, 66.27% of respondents who identified an email as definitely phishing opted for move to spam as their action. However, the percentage of participants who rated the email as genuine (definitely genuine + probably genuine) did not closely match the percentage who chose to click link. Interestingly, for emails generated by WebSearch agents, the click rate was much higher than the rate of perceived genuineness, a pattern not observed with vanilla LLM-generated emails. This discrepancy was driven by participants marked as unsure or even phishing; 13.34% of participants stated that they would clicked on links in WebSearch agent-generated emails despite doubting their authenticity. (For vanilla LLM-generated emails, 10% and 3.34% did so for Claude and GPT, respectively.) Table 11 (Appendix E.3) presents a contingency table showing the percentage distribution of participants’ actions (Q5) relative to their perception of authenticity (Q6). These results highlight that the WebSearch agents not only provided a sense of credibility but also effectively utilized target information from the web to engage the recipients’ interest, leading to higher link-clicking rates.
@@
Content Analysis. We analyzed which email aspects influence participants’ perceptions of authenticity using chi-squared tests. From the analysis, the detail of email content and the accuracy of those details in aligning with the target’s information played statistically significant roles in participants’ assessment of an email’s authenticity. Table 8 in Appendix E.2 provides statistical results of the chi-squared tests.
Building on these statistical results, we further examined the influence of web accessibility on the two factors. Figure 11 presents a comparative analysis of the percentage of phishing emails that incorporate specific content details and how well this content aligns with the targets’ actual information, based on participants’ evaluations. When comparing GPT and Claude, it was noted that GPT-generated emails were less detailed, regardless of web usage. We also observed that emails crafted by WebSearch agents were significantly richer in detail and more accurately reflected the target’s personal data than those generated by vanilla LLMs. This difference was particularly pronounced in the category of specific expertise. According to the results of our chi-squared tests in Table 9 (Appendix E.2), the alignment of content with the target’s specific expertise had the most significant impact on participants’ perception of the email’s authenticity. By leveraging detailed information about the target’s expertise from the web, WebSearch agents were able to craft more convincing and effective phishing emails.
@@ 7 Approachability of Cyberattacks
In this section, we examine the practicality of using LLM agents for targeted cyberattacks by assessing two primary factors: cost and safeguard capability. Understanding these elements is crucial to determining the practicality of employing LLM agents for malicious purposes in real-world scenarios. The time and financial costs associated with the attacks are measured, followed by an assessment of bypassing LLM safeguards.
@@ 7.1 Cost Analysis
Conducting attacks incurs both time and financial costs, which can highly impact the practicality of using LLM agents for cyberattacks. We examine how quickly and at what cost attackers can carry out cyberattacks using commercially available LLMs. We measured the time spent and the number of tokens consumed during each attack scenario using WebSearch agent for each individual target. For a cost analysis of the WebNav agent, please refer to Appendix D.1. To simplify our calculations, we considered token usage as the sum of input and output tokens, applying the output token rate, which is typically more expensive than the input token rate. This approach ensures we estimate the upper limit of API costs. For the models utilized, the costs are as follows: GPT 4o [41] and Claude 1.5 Sonnet [5] at \$15/1M tokens, and Gemini-1.5 Flash [20] at \$0.30/1M tokens. To measure the number of tokens, we utilized the tiktoken tokenizer [42] for GPT, as specified by OpenAI [42]. For Claude and Gemini, we used the token counting feature provided by their APIs [3, 13].
On average, collecting PII for an individual  [^13] We sample 50 professors identified in Section 4. required 9.1 seconds and incurred a cost of 2.2 cents (GPT: 6.7s, 2.1¢; Claude: 11.4s, 2.2¢). For generating impersonation posts, GPT, Claude, and Gemini took an average of 9.4 seconds, 21.7 seconds, and 3.6 seconds, respectively, with costs under 3 cents (GPT: 2.2¢; Claude: 3.0¢; Gemini: 0.03¢). For creating phishing emails, GPT and Claude took an average of 7.1 seconds and 22.7 seconds, respectively, with costs below 4 cents (GPT: 2.7¢; Claude: 3.9¢). See Appendix D for the cost statistics.
@@ 7.2 Safeguard Capability
Spear Phishing Email Generation. The assessment of safeguard in LLMs for spear phishing email generation involved varying two key factors: 1) the purpose of the email (general, request login credentials, request money), and 2) the type of email address used (institutional or Gmail). We generated emails using the email addresses of the authors, with their consent.
Gemini shows the strongest safeguards by refusing to generate emails in all scenarios of vanilla LLM and WebSearch agent settings (Table 7 in Appendix D.3), responding with messages like “I cannot fulfill your request. Creating emails without consent is unethical and potentially illegal”. All models without web accessibility refuse generating content when the email’s purpose involves requesting login credentials or money, indicating these topics trigger stricter safeguard mechanisms. Conversely, general-purpose emails are less likely to be blocked, with GPT and Claude generating responses in these cases. GPT and Claude are more likely to generate an email when provided with an institutional email address, compared to a Gmail address.

### v3

@@ 5.1 Experimental Setup
Evaluation. We examined the impact of web accessibility on impersonation by conducting an A/B test. Impersonation posts were generated using three different methods: (1) vanilla LLMs, (2) WebSearch agents, and (3) WebNav agents. We then conducted pairwise comparisons to determine which post was more likely to have been written by the target individual. As evaluators, we employed LLMs, which are recognized for their high accuracy in tasks such as fake news detection and machine-generated text detection, even without additional fine-tuning [53, 14].
We enabled the web search function to enhance the accuracy of the LLM evaluations by leveraging internet-available information. Evaluators were presented with posts from two different models (e.g., vanilla LLM vs. WebSearch agent, or WebSearch agent vs. WebNav agent) and indicated which post more accurately impersonated the target individual, or marked ‘unsure’ if the posts were similarly effective. We randomize the placement of texts (A or B) to mitigate order bias. Furthermore, to minimize potential model bias, each pair of posts was evaluated by three different models: GPT, Claude, and Gemini. The final outcome was determined based on the majority vote from three LLM evaluations.
In addition, to determine how likely the impersonation posts were perceived to be authentic, we conducted a Yes/No test. Similar to the A/B test, we utilized the three LLM evaluators and to assess whether the texts appeared to have been authored by the target individuals. The templates for the A/B test and Yes/No test are provided in Appendix C.
Validation of the LLM-judging-LLM output. We assessed the reliability of our LLM evaluators by comparing them with human reviewers. We evaluated a total of 270 posts in our validation process. This number was derived from the formula 270 = 15 (number of professors) × 2 (number of claims) × 3 (number of language models: GPT, Claude, Gemini) × 3 (number of settings: LLM, WebSearch agent, WebNav agent). We sampled 15 professors, each represented in 18 posts generated across the various configurations. The annotation guidelines provided to human reviewers were the same as those presented as prompts to the LLM evaluators.
We compared LLM evaluations with human evaluations to assess the models’ accuracy against human judgment. For the A/B Test, the average agreement rate between humans and LLM evaluators was 92.8% for the first claim and 85.0% for the second claim. For the Yes/No Test, these rates were 92.6% and 91.1%, respectively.
@@ 5.2 Main Results
Figure 8 shows an impersonation post of a well-known figure, generated by the agent with GPT. [^3] For visualization, we prompted the WebSearch agent to generate a tweet of up to 300 words with the second claim. We provide the agent with only the name information; however, it synthesizes information about his expertise, alma mater, and companies to craft a tweet that convincingly appears to be authored by him. The proceeding tests are conducted using target professors according to our attack pipeline.
Comparison Based on Tool Usage. We conducted comparative analysis of impersonation posts generated by WebNav agents and WebSearch agents, followed by comparisons between WebSearch agents and vanilla LLMs, to examine their effectiveness based on their capability. Table 2 shows the proportion of choices made in an A/B test. Cases where the LLMs or agents refused to generate content were excluded from our calculations (these will be discussed later in this section). The results indicate that the effectiveness of impersonation increases with access to additional tools across all models. Specifically, WebNav agents were more effective than WebSearch agents, and WebSearch agents outperformed vanilla LLMs. The differences in effectiveness were more pronounced when comparing WebSearch agents with vanilla LLMs, suggesting that adding search capabilities (WebSearch agents) provides a significant boost in performance for tasks requiring impersonation abilities.
Table 2: A/B test results. The percentages represent each decision relative to the total number of decisions made. WebN and WebS represent WebNav and WebSearch, respectively. claim1 refers to “Recommend researching AI” and claim2 to “LLM is highly secure against potential misuse”.
@@
Table 3: Yes/No test results. Percentages represent the proportion of “Yes” responses.
@@
Gen WebN Agent WebS Agent LLM WebN Agent WebS Agent LLM
@@
Authenticity Evaluation. We examine how the texts created by each of the agents and vanilla LLMs are regarded in terms of authenticity. Table 3 shows the percentage of “yes” responses (deemed as authentic) by the LLM evaluator for posts created by vanilla LLMs and LLM agents. The results are similar to those observed in the A/B test: agent-generated posts had a higher “yes” response rate than LLM-generated posts. WebNav agents performed the best; the GPT and Claude agents achieved success rates of up to 93.9% and 85.7%, respectively. Compare to GPT and Claude, Gemini showed poor performance in most cases.
@@
6 Spear Phishing Email Generation
In this section, we investigate wthe capability of LLM agents to autonomously generate personalized phishing emails using only an email address, without additional human intervention.
@@ 6.1 Experimental Setup
We conducted two pilot studies, which showed similar results for WebNav and WebSearch Agents. Due to cost and time constraints, we excluded WebNav-generated emails from the main evaluation.
@@
We categorized participants into two groups: academic researchers and non-academic professionals. In Section 4, we noted that information about academic researchers is publicly available online, facilitating easy collection by attackers. Consequently, we assumed that academic researchers would be more susceptible to targeted phishing attacks, prompting their inclusion in our study. We also included non-academic professionals to examine phishing susceptibility across diverse organizational contexts, expanding the analysis beyond academic settings. We defined non-academic professionals as individuals employed within non-academic organizations.
We recruited academic researchers by posting experiment descriptions within university research communities. Each community consists of people engaged in research including professors, graduate students, and undergraduate students. We recruited non-academic professionals through industry practitioners who have previously collaborated with the authors. In total, we recruited 60 participants through Google Forms and categorized them based on the provided institutional email addresses to analyze response variations across institutional types: 55% were academic researchers and 45% were non-academic professionals. All participants were informed about an incentive of approximately \$7.50 for their participation.
We emphasize that participants did not directly receive these emails; instead, their responses were assessed through a Google survey. [^4] The process with ethical/privacy-preserving measures is explained in Appendix E.1. The rationale for adopting this survey format over a simulated phishing attack will be discussed in Section 8. Each participant evaluated seven phishing emails generated using their own email as a target. The types of emails will be described in Email Design part.
@@
Questionnaire Design. We designed seven questions to build an understanding of how participants perceive and interact with the emails, from surface content to deeper implications of authenticity and expected actions. The first two questions focus on the content of the emails, asking participants to identify the presented information (Q1) and to evaluate its accuracy against their actual personal details (Q2). Subsequent questions explore perceptions of the sender and recipient, including whether the recipient identified the sender (Q3) and if the recipient’s name correctly matched their real name (Q4). We then inquire about the actions participants would likely take upon receiving the email (Q5). The last question investigates perceived authenticity of the email (Q6), asking participants to rate how genuine or fraudulent it appeared and to identify specific elements that influenced their judgment (Q7). The questionnaire is available at this link [^5] https://zenodo.org/doi/10.5281/zenodo.13691327.
@@ 6.2 Main Results
We first analyze the effectiveness and content of general-purpose emails created by vanilla LLMs and WebSearch agents. Then, we examine the results based on the participants’ groups. We also analyze the efficacy of emails in relation to their intended purpose and the sender’s institution. In Tables 7 (Appendix E.3), we report the comprehensive survey results concerning participants’ anticipated actions for various email types.
@@ 6.2.1 Overall Analysis
Effectiveness Analysis. Figure 11 (a) presents the survey results of actions participants would likely take upon receiving the email (Q5). Our primary focus was on the proportion of participants clicking on links, directly aligning with the attacker’s goal. The results indicate that WebSearch agents prompt more link clicks than vanilla LLMs, with the WebSearch agent using Claude achieving the highest click rate at 26.67%. [^6] Expert-crafted spear phishing emails in [11] show 26.6% click rate. Detailed discussion in Section 6.2.3. In the case of GPT, the WebSearch agent’s click rate was twice that of the vanilla LLM. When comparing Claude and GPT, Claude was more effective, with its vanilla LLM variant achieving a slightly higher click rate than the WebSearch agent using GPT. The disparities in effectiveness between Claude and GPT may be attributed to differences in email content, which will be explored further in the subsequent section.
When comparing “actions” (Q5) and “perceived authenticity” (Q6) in Figure 11, the percentage of participants identifying the email as definitely phishing closely aligned with those opting move to spam. On average, 66.27% of respondents who identified an email as definitely phishing opted for move to spam as their action. However, the percentage of participants who rated the email as genuine (definitely genuine + probably genuine) did not closely match the percentage who chose to click link. Interestingly, for emails generated by WebSearch agents, the click rate was much higher than the rate of perceived genuineness, a pattern not observed with vanilla LLM-generated emails. This discrepancy was driven by participants marked as unsure or even phishing; 13.34% of participants stated that they would click on links in WebSearch agent-generated emails despite doubting their authenticity. In contrast, for vanilla LLM-generated emails, 3.34% and 10% did so for GPT and Claude, respectively. Table 8 (Appendix E.4) presents a contingency table showing the percentage distribution of participants’ actions (Q5) relative to their perception of authenticity (Q6). These results highlight that the WebSearch agents not only provided a sense of credibility but also effectively utilized target information from the web to engage the recipients’ interest, leading to higher link-clicking rates.
@@
Content Analysis. We analyzed which email aspects influence participants’ perceptions of authenticity using chi-squared tests. From the analysis, the alignment of the details of email content and the target’s personal information played statistically significant roles in participants’ assessment of an email’s authenticity. Table 6 in Appendix E.2 provides statistical results of the chi-squared tests.
Building on these statistical results, we further examined the influence of web accessibility on the two factors. Figure 12 presents a comparative analysis of the percentage of phishing emails that incorporate specific content details and how well this content aligns with the targets’ actual information, based on participants’ evaluations. When comparing GPT and Claude, it was noted that GPT-generated emails were less detailed, regardless of web usage. We also observed that emails crafted by WebSearch agents were significantly richer in detail and more accurately reflected the target’s personal data than those generated by vanilla LLMs. This difference was particularly pronounced in the category of specific expertise. The results from our chi-squared tests in Table 6 (Appendix E.2), indicate that the alignment of content with a target’s specific expertise significantly influenced the perceived authenticity of the emails. By leveraging detailed web-sourced information about the target’s expertise, WebSearch agents were able to craft more convincing and effective phishing emails.
@@ 7 Approachability of Cyberattacks
In this section, we examine the practicality of using LLM agents for targeted cyberattacks by assessing two primary factors: cost and safeguard capability. Understanding these elements is crucial to determining the practicality of employing LLM agents for malicious purposes in real-world scenarios.
@@ 7.1 Cost Analysis
Conducting attacks incurs both time and financial costs, which can highly impact the practicality of using LLM agents for cyberattacks. We examine how quickly and at what cost attackers can carry out cyberattacks using commercially available LLMs. We measured the time spent and the number of tokens consumed during each attack scenario using WebSearch agent for each individual target. To simplify our calculations, we considered token usage as the sum of input and output tokens, applying the output token rate, which is typically more expensive than the input token rate. This approach ensures we estimate the upper limit of API costs. For the models utilized, the costs are as follows: GPT 4o [49] and Claude 1.5 Sonnet [7] at \$15/1M tokens, and Gemini-1.5 Flash [25] at \$0.30/1M tokens. To measure the number of tokens, we utilized the tiktoken tokenizer [50] for GPT, as specified by OpenAI [50]. For Claude and Gemini, we used the token counting feature provided by their APIs [5, 16].
On average, collecting PII for an individual required 9.1 seconds and incurred a cost of 2.2 cents (GPT: 6.7s, 2.1¢; Claude: 11.4s, 2.2¢). For generating impersonation posts, GPT, Claude, and Gemini took an average of 9.4, 21.7, and 3.6 seconds, respectively, with costs under 3 cents (GPT: 2.2¢; Claude: 3.0¢; Gemini: 0.03¢). For creating phishing emails, GPT and Claude took an average of 7.1 and 22.7 seconds, respectively, with costs below 4 cents (GPT: 2.7¢; Claude: 3.9¢). See Appendix D for the cost statistics.
@@ 7.2 Safeguard Capability
Spear Phishing Email Generation. The safeguard assessment in LLMs against spear phishing email generation involved varying two key factors: 1) the purpose of the email (general, request login credentials, request money), and 2) the type of email address used (institutional or Gmail). We generated emails using the email addresses of the authors, with their consent.
Gemini shows the strongest safeguards by refusing to generate emails in all scenarios of vanilla LLM and WebSearch agent settings, responding with messages like “I cannot fulfill your request. Creating emails without consent is unethical and potentially illegal”. All models without web accessibility refuse generating content when the email’s purpose involves requesting login credentials or money, indicating these topics trigger stricter safeguard mechanisms. Conversely, general-purpose emails are less likely to be blocked, with GPT and Claude generating responses in these cases. GPT and Claude are more likely to generate an email when provided with an institutional email address, compared to a Gmail address.
@@
8 Discussion and Limitations
8.1 Adoption of Survey Format
In our study, we chose a survey-based approach to understand user behavior towards phishing emails due to the following considerations.
Understanding Intentions and Reasoning. Survey-based research allows us to explore the reasoning and intentions behind participants’ actions, providing insights that are difficult to capture through mere analysis of click rates. Such methods have been endorsed in prior research [61, 31, 39] as effective means of understanding user behavior in security settings.
Addressing Ethical Concerns. Simulating phishing attacks often involve deceptive practices that raise significant ethical concerns [59, 13, 17]. Prior studies using simulations tend to be restricted to single organizations where comprehensive controls and institutional reviews are possible [37, 11]. These controlled environments, while beneficial for managing ethical risks, restrict the diversity of the participant pool. By opting for a survey-based approach, we were able to gather valuable data across 25 different organizations without encountering ethical risks.
8.2 Defense
We suggest defense strategies against the unauthorized collection of PII by LLMs utilizing web-based tools. First, LLM service vendors can implement a rule requiring LLMs that use web crawling tools to check and adhere to the directives specified in the robots.txt files of websites. These files serve as an international recommendation that either allows or restricts web crawlers from collecting site and webpage data. By respecting these directives, LLMs can prevent the inadvertent or intentional scraping of sensitive information from restricted areas of websites. Additionally, scalable safeguards can be deployed, as further discussed in Section 8.3.
Second, PII providers (e.g., website managers) responsible for handling sensitive information should proactively control and reduce its online exposure. A straightforward measure involves crafting and regularly updating robust robots.txt files that explicitly block crawlers from accessing pages containing personal information. In addition, they can employ more creative tactics to “trick” LLMs into gathering incorrect or incomplete PII. For instance, a website might display scrambled or false personal data to automated crawlers, while showing the correct information only after a user takes another intentional action. This approach ensures that human visitors see the accurate details, but an LLM performing automated scraping inadvertently collects erroneous information, thereby reducing the potential for unauthorized access or misuse.
8.3 Risks with Expanding Capabilities
The evolution of LLMs has not only linked their expanding capabilities to increased proficiency in executing complex tasks but also highlighted their potential misuse in scenarios such as cyberattacks. Our findings demonstrate that as LLMs are enabled with additional tools, their effectiveness in cyberattacks is significantly enhanced. This correlation between enhanced capabilities and the associated risks underscores the necessity for proportionately stronger safeguards.
In response to these growing concerns, companies like Anthropic have developed scalable safeguards with the potential risks [3]. This approach involves setting specific thresholds at which the abilities of an AI system necessitate enhanced safeguards. Upon reaching these thresholds, the system automatically enforces stronger security measures tailored to the level of risk posed by the AI’s capabilities. As the capabilities of LLMs increase, implementing scalable and adaptive security measures will be paramount to ensure their safe and ethical utilization.
8.4 Limitations
In Section 5, one of our primary objectives was to evaluate the effectiveness of LLM agents in generating impersonation posts according to their capability. Understanding the link between tool usage and impersonation effectiveness required experiments to be conducted under uniform conditions, such as consistent target claims. However, collecting authentic posts from our targets—CS professors— regarding specific claims was infeasible. Thus, we were unable to use actual samples for comparison. Despite this, the inability to use authentic posts did not undermine our study’s primary objective.
We also relied on LLM evaluators to judge the impersonation posts. Although these evaluators showed high agreement rates compared to human reviewers, they may not fully capture the nuanced reasoning and contextual awareness that human experts bring.
In Section 6, we focused on analyzing how participants interacted with LLM-generated phishing emails. While including synthetic non-phishing emails could provide a valuable baseline for comparison, this aspect was not incorporated into the current study. We suggest it as a promising direction for future research to enhance our understanding of user responses to both genuine and phishing emails.